# EgoDex: Learning Dexterous Manipulation from Large-Scale Egocentric Video

### v1

@@ Abstract
Imitation learning for manipulation has a well-known data scarcity problem. Unlike natural language and 2D computer vision, there is no Internet-scale corpus of data for dexterous manipulation. One appealing option is egocentric human video, a passively scalable data source. However, existing large-scale datasets such as Ego4D do not have native hand pose annotations and do not focus on object manipulation. To this end, we use Apple Vision Pro to collect EgoDex: the largest and most diverse dataset of dexterous human manipulation to date. EgoDex has 829 hours of egocentric video with paired 3D hand and finger tracking data collected at the time of recording, where multiple calibrated cameras and on-device SLAM can be used to precisely track the pose of every joint of each hand. The dataset covers a wide range of diverse manipulation behaviors with everyday household objects in 194 different tabletop tasks ranging from tying shoelaces to folding laundry. Furthermore, we train and systematically evaluate imitation learning policies for hand trajectory prediction on the dataset, introducing metrics and benchmarks for measuring progress in this increasingly important area. By releasing this large-scale dataset, we hope to push the frontier of robotics, computer vision, and foundation models.
@@ 2.3 Learning from Human Video
Most similar to our work is EgoMimic [15], which proposes the collection of egocentric video and paired 3D hand tracking. The primary difference is scale: while EgoMimic collects around 4 hours of data, we collect 829 hours with a much broader data and task distribution. We also collect more dexterous annotations, critical for downstream manipulation: 3D positions and orientations for the upper body including the head, shoulders, arms, and 25 joints in each hand, whereas EgoMimic collects only the wrist positions. Lastly, a preview of EgoDex appeared in PH2D [30], which used a 3% subset of EgoDex for successful human-to-robot transfer.
@@ 3.2 Data Collection
To streamline data collection, data is recorded in sessions: approximately 10-15 minute segments that consist of many individual episodes, where episode boundaries are indicated by a “pause” and subsequent “resume” of recording from the data collection app. Raw video is compressed to facilitate data transfer, upload, download, and storage. Without the use of modern video compression algorithms, the raw data would take over 500 TB of disk space, about 250$\times$ its current size. At training time, data is loaded efficiently with PyTorch torchcodec [37], which only decodes the desired frames in the sampled batch of data.
@@ 3.5 Diversity
Prior works [16, 26] identify several potential axes of demonstration diversity: viewpoint diversity, task diversity, scene diversity, object diversity, and more. In EgoDex, the emphasis is diversity in dexterous manipulation behaviors. Tasks and objects vary such that the required dexterity ranges far beyond pick-and-place, the primary behavior in most robot teleoperation datasets. For example, tasks include tightening a screw, tying shoelaces, dealing cards, flipping pages, catching tennis balls, and slotting batteries. The task distribution covers a wide range of everyday household manipulation tasks that can be performed on a tabletop surface. There is also a significant amount of basic pick-and-place with diverse objects as well as the benchmark tasks from the FurnitureBench assembly benchmark [13]. The full list of 194 tasks provided in Appendix  A.1.
@@ 4.1 Action Representation
Given the full set of skeletal joints in the EgoDex dataset, many action representations are possible: wrist positions, wrist orientation, positions of fingertips, and so on. Since we focus on dexterous manipulation, we choose a representation that both captures sufficient bimanual dexterity. Specifically, the action $\mathbf{a_{t}}$ at time $t$ is represented as the 3D position of each wrist, the 6D orientation of each wrist (as suggested by [43]), and the 3D position of each fingertip. Thus, each action has a total dimensionality of 2 hands $\times$ (3 + 6 + (3 $\times$ 5 fingertips)) = 48. In practice, actions are predicted in chunks over a fixed time horizon. All poses are represented in the camera frame.
@@ 4.3 Evaluation Metrics
Thus, for each benchmark task we evaluate performance with a “best of $K$” metric. For each data point in the full test set, we sample the trained model $K$ times to capture different possible modes. We then compute the distance between the ground truth trajectory and the trajectory closest to it out of the $K$ samples, where “distance” is calculated as the Euclidean distance between predicted 3D keypoint positions and their ground truth 3D counterparts, averaged over each timestep in the predicted chunk and each of the 12 keypoints (i.e., the wrist and fingertips of each hand). Intuitively, this value can be interpreted as the average positional error in 3D space between ground truth and prediction in meters. The final value is averaged over the full test set. For deterministic models, the value is the same regardless of $K$; for stochastic models, the value improves as $K$ increases, as the model gets more chances to sample the ground truth mode.
@@ 6 Research Use Cases
Given a hypothetical robot embodiment that mimics the visual appearance, kinematics, and dynamics of human arms and hands, a control policy mapping egocentric video to 3D arm and hand pose deltas such as those trained in Section 5 would be deployable zero-shot without any additional data or fine-tuning. Such an embodiment would lie within the training data distribution, both visually and dynamically.
@@
EgoDex can be used for learning tasks such as action recognition and human-object interaction detection from egocentric videos. Datasets like EPIC-KITCHENS [8] has demonstrated the value of egocentric video for recognizing and anticipating daily actions, and even more challenging tasks like detecting active objects and predicting state changes from egocentric video. Researchers can also study which objects are involved in each action and how. For example, modeling the contact points, grasps, and trajectories when using a tool (screwdriver, scissors, etc.). A related task is learning object affordances, i.e., understanding what actions each object supports.
@@
Recent advances in large-scale diffusion models have significantly enhanced the capabilities of language-conditioned video generation, producing temporally consistent and semantically accurate visual narratives from natural language inputs [18, 29, 24]. These generative frameworks have demonstrated potential not only in creating realistic and detailed video content but also as world models for decision-making tasks, supporting reinforcement learning agents by simulating future outcomes based on predicted visual dynamics [40, 3, 41, 12]. Despite these impressive advancements, there remains a substantial research gap in video generation and world modeling from an egocentric viewpoint. Egocentric perspectives present unique challenges, including managing significant viewpoint variability, maintaining temporal and spatial consistency amid frequent camera movements, and accurately reflecting agent-centric interactions and intentions. Since EgoDex provides large-scale video data with 3D pose and language annotations, it enables us to train an egocentric foundation world model.
@@ A.3 Training Details
Each model is trained and evaluated on a single node with 96 logical CPUs (48 physical CPUs) and 8 NVIDIA A100 GPUs each with 80GB RAM. Training is run for 50,000 gradient steps with a batch size of 2048 (256 per GPU with data parallelism), at which point training and validation loss plateau. The full training run takes approximately 72 hours. The models are optimized with Adam and a learning rate of 1e-4. Image observations are resized to 224 $\times$ 224 and sent through a pretrained ResNet encoder, while language annotations are passed through a frozen CLIP [31] encoder. DDPM and FM models are trained and evaluated with 16 sampling steps. All other hyperparameters are the defaults from the X-IL codebase [14].

### v2

@@ Abstract
Imitation learning for manipulation has a well-known data scarcity problem. Unlike natural language and 2D computer vision, there is no Internet-scale corpus of data for dexterous manipulation. One appealing option is egocentric human video, a passively scalable data source. However, existing large-scale datasets such as Ego4D do not have native hand pose annotations and do not focus on object manipulation. To this end, we use Apple Vision Pro to collect EgoDex: the largest and most diverse dataset of dexterous human manipulation to date. EgoDex has 829 hours of egocentric video with paired 3D hand and finger tracking data collected at the time of recording, where multiple calibrated cameras and on-device SLAM can be used to precisely track the pose of every joint of each hand. The dataset covers a wide range of diverse manipulation behaviors with everyday household objects in 194 different tabletop tasks ranging from tying shoelaces to folding laundry. Furthermore, we train and systematically evaluate imitation learning policies for hand trajectory prediction on the dataset, introducing metrics and benchmarks for measuring progress in this increasingly important area. By releasing this large-scale dataset, we hope to push the frontier of robotics, computer vision, and foundation models. EgoDex is publicly available for download at https://github.com/apple/ml-egodex.
@@ 2.3 Learning from Human Video
Most similar to our work is EgoMimic [15], which proposes the collection of egocentric video and paired 3D hand tracking. The primary difference is scale: while EgoMimic collects around 4 hours of data, we collect 829 hours with a much broader data and task distribution. We also collect more dexterous annotations, critical for downstream manipulation: 3D positions and orientations for the upper body including the head, shoulders, arms, and 25 joints in each hand, whereas EgoMimic collects only the wrist positions. Lastly, a preview of EgoDex appeared in Humanoid Policy $\sim$ Human Policy [30], which used a 3% subset of EgoDex for successful human-to-robot transfer.
@@ 3.2 Data Collection
[…]
@@ 3.5 Diversity
Prior works [16, 26] identify several potential axes of demonstration diversity: viewpoint diversity, task diversity, scene diversity, object diversity, and more. In EgoDex, the emphasis is diversity in dexterous manipulation behaviors. Tasks and objects vary such that the required dexterity ranges far beyond pick-and-place, the primary behavior in most robot teleoperation datasets. For example, tasks include tightening a screw, tying shoelaces, dealing cards, flipping pages, catching tennis balls, and slotting batteries. The task distribution covers a wide range of everyday household manipulation tasks that can be performed on a tabletop surface. There is also a significant amount of basic pick-and-place with diverse objects as well as the benchmark tasks from the FurnitureBench assembly benchmark [13]. The full list of 194 tasks is provided in Appendix  A.1.
@@ 4.1 Action Representation
Given the full set of skeletal joints in the EgoDex dataset, many action representations are possible: wrist positions, wrist orientation, positions of fingertips, and so on. Since we focus on dexterous manipulation, we choose a representation that captures sufficient bimanual dexterity. Specifically, the action $\mathbf{a_{t}}$ at time $t$ is represented as the 3D position of each wrist, the 6D orientation of each wrist (as suggested by [43]), and the 3D position of each fingertip. Thus, each action has a total dimensionality of 2 hands $\times$ (3 + 6 + (3 $\times$ 5 fingertips)) = 48. In practice, actions are predicted in chunks over a fixed time horizon. Poses are expressed in the current camera frame [15], and each action chunk is a relative trajectory [7].
@@ 4.3 Evaluation Metrics
Thus, for each benchmark task we evaluate performance with a “best of KK” metric. For each data point in the full test set, we sample the trained model $K$ times to capture different possible modes. We then compute the distance between the ground truth trajectory and the trajectory closest to it out of the $K$ samples, where “distance” is calculated as the Euclidean distance between predicted 3D keypoint positions and their ground truth 3D counterparts, averaged over each timestep in the predicted chunk and each of the 12 keypoints (i.e., the wrist and fingertips of each hand). Intuitively, this value can be interpreted as the average positional error in 3D space between ground truth and prediction in meters. The final value is averaged over the full test set. For deterministic models, the value is the same regardless of $K$; for stochastic models, the value improves as $K$ increases, as the model gets more chances to sample the ground truth mode.
@@ 6 Research Use Cases
Given a hypothetical robot embodiment that mimics the visual appearance, kinematics, and dynamics of human arms and hands, a control policy mapping egocentric video to 3D arm and hand pose deltas such as those trained in Section 5 may be deployable zero-shot without additional data or fine-tuning. Such an embodiment would lie within the training data distribution, both visually and dynamically.
@@
EgoDex can be used for learning tasks such as action recognition and human-object interaction detection from egocentric videos. Datasets like EPIC-KITCHENS [8] has demonstrated the value of egocentric video for recognizing and anticipating daily actions, and even more challenging tasks like detecting active objects and predicting state changes from egocentric video. Researchers can also study which objects are involved in each action and how. For example, one could model the contact points, grasps, and trajectories when using a tool (screwdriver, scissors, etc.). A related task is learning object affordances, i.e., understanding what actions each object supports.
@@
Recent advances in large-scale diffusion models have significantly enhanced the capabilities of language-conditioned video generation, producing temporally consistent and semantically accurate visual narratives from natural language inputs [18, 29, 24]. These generative frameworks have demonstrated potential not only in creating realistic and detailed video content but also as world models for decision-making tasks, supporting reinforcement learning agents by simulating future outcomes based on predicted visual dynamics [40, 3, 41, 12]. Despite these impressive advancements, there remains a substantial research gap in video generation and world modeling from an egocentric viewpoint. Egocentric perspectives present unique challenges, including managing significant viewpoint variability, maintaining temporal and spatial consistency amid frequent camera movements, and accurately reflecting agent-centric interactions and intentions. Since EgoDex provides large-scale video data with 3D pose and language annotations, it enables the training of an egocentric world model.
@@ A.3 Training Details
Each model is trained and evaluated on a single node with 96 logical CPUs (48 physical CPUs) and 8 NVIDIA A100 GPUs each with 80GB RAM. Training is run for 50,000 gradient steps with a batch size of 2048 (256 per GPU with data parallelism), at which point training and validation loss plateau. The full training run takes approximately 72 hours. The models are optimized with Adam and a learning rate of 1e-4. Image observations are resized to 224 $\times$ 224 and sent through a pretrained ResNet encoder, while language annotations are passed through a frozen CLIP [31] encoder. Only the current image observation and proprioceptive state are passed as input to the policy (i.e., no history); adding history may improve performance. DDPM and FM models are trained and evaluated with 16 sampling steps. All other hyperparameters are the defaults from the X-IL codebase [14].
@@ A.4 Dataset Access
The full dataset as well as sample Python code for data loading are available at https://github.com/apple/ml-egodex. Instructions for downloading the dataset are in the README.md, and repeated below for convenience.